OpenStudy (anonymous):

Help me solve this radical equation! Its easy at first except for some odd reason i end up with a not right answer :/ Please help me :)

OpenStudy (anonymous):

\[\sqrt[3]{5x-1}-1=2\]

OpenStudy (anonymous):

I think I'm messing up with my negative and positive signs

OpenStudy (amistre64):

N-1 = 2 , N = 3 3rt(M) = 3, M= 27 5x-1 = 27, x = 28/5 maybe

OpenStudy (anonymous):

ill check

OpenStudy (anonymous):

*hug* Thank you :D